What is the pH of an aqueous solution of 3.99×10-2 M nitric acid


pH = 

1
Expert's answer
2022-05-09T14:57:05-0400

Solution:

nitric acid = HNO3


HNO3 is a strong acid because it dissociates almost completely.

HNO3(aq) = H+(aq) + NO3(aq)

Therefore,

[H+] = C(HNO3) = 3.99×10−2 M


We can convert between [H+] and pH using the following equations:

pH = −log[H+]

pH = −log(3.99×10−2) = 1.399 = 1.4

pH = 1.4


Answer: pH = 1.4